# Content Accuracy and Quality Aware Resource Allocation Based on LP-Guided DRL for ISAC-Driven AIGC Networks

### v1

@@ 4.4 Complexity Analysis
For the overall complexity of LPDRL-F, if standard LP solvers, such as simplex methods, are employed to solve ComRA sub-problem, the complexity of simplex methods cannot be ignored, and the overall complexity of LPDRL-F would be $O\left({{2^{K}}+\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}}\right)$. In contrast, the time complexity of our proposed LPDRL-F algorithm can be summarized as $O\left({\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}}\right)$. This is because existing DRL typically use 32-256 hidden layer neurons [36]. According to 3GPP standards [37, 38], the number of immersive VR/XR interaction users $K$ is at the order of 10. Thus, $K\log K\ll\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}$, and the complexity of the RCE algorithm can be neglected.
5 Performance Evaluation
@@
Minimum CAQA $\eta_{k,\text{min}}$ $\sim U[0.4,0.45]$
@@
Simulation parameters of the ISAC-based AIGC environment and LPDRL-F are listed in Table I and Table II, respectively [17, 18, 20, 21, 39, 40]. Large-scale path loss is given by $128.1+37.6{\rm{lo}}{{\rm{g}}_{10}}\left({d/1000}\right)$, where $d$ is the distance between the ISAC device and the user in meters [31]. $\delta^{2}$} is -174 dBm/Hz. The simulation consists of 6000 training episodes, each with 50 iterations, and 10000 testing episodes with 10 iterations. To fully evaluate the algorithm with diverse scenarios, user positions are randomly reset at the beginning of each episode, and users’ task priority level and CAQA requirements are reset at each time slot.
@@
  • Sensing accuracy and quality aware with fixed generating step (SAQA-FG): Consider the impact of sensing accuracy and communication quality while assuming perfect generation without error. i.e., AEG=0. Sensing energy allocation is solved using DRL-F with an action filter, and ComRA is optimized by LP. Since AEG is not considered, the generating steps are fixed to a given value. Comparing SAQA-FG with CAQA-AIGC can show the importance of considering AEG in ISAC-AIGC networks.
@@
Then, Fig. 7 shows the AvgCAQA performance as a function of the maximum total energy $E_{\max}$}. It can be seen that for all investigated algorithms, AvgCAQA improves as $E_{\max}$} increases, due to more available energy for both sensing and communication. The proposed CAQA-AIGC with LPDRL-F consistently achieves the highest AvgCAQA. Similar to Fig. 6, this is because with the assumption of perfect content accuracy, CGQ-FSG-$\alpha$ is designed to ensure a fixed sensing energy proportion and then allocate the remaining energy for communication based on a CGQ-oriented strategy. In a practical system with imperfect content accuracy, when the total energy is small, allocating a larger proportion of sensing energy can achieve higher content accuracy (see Fig. 8). However, it also results in lower content quality since less resource is available for communications (see Fig. 9). Compared to CGQ-FSG-$\alpha$, CAQA-AIGC with LPDRL-F can achieve a better balance between the content accuracy and quality and outperforms CGQ-FSG-$\alpha$ when $E_{\max}$} is small. As $E_{\max}$} increases, CGQ-FSG-$\alpha$ can allocate sufficient energy for communication and enhance content quality (see Fig. 9). However, since a fixed sensing energy is employed by CGQ-FSG-$\alpha$, it cannot fully exploit the advantage of high total energy to further enhance content accuracy. Different to CGQ-FSG-$\alpha$, CAQA-AIGC is aware of the content accuracy in a practical system. It can optimize the sensing-generating-communication resource allocation to fully exploit the high total energy. As shown in Fig. 8 and Fig. 9, when $E_{\max}$} is large, CAQA-AIGC achieves a notably better content accuracy with a slightly worse content quality. Overall, the AvgCAQA performance of CAQA-AIGC is always better than that of CGQ-FSG-$\alpha$.
@@
To evaluate the impact of AEG (one of the sources affecting content accuracy) and the effectiveness of adaptive generating step optimization, Fig. 10 shows AvgCAQA as a function of the generating steps with different computing capacities. With CAQA-AIGC, the generation step is jointly optimized with sensing and communication resources, yielding consistently high AvgCAQA across all computation capacities. In contrast, SAQA-FG uses fixed generation steps. As the number of generating steps increases, AvgCAQA initially improves due to higher generating accuracy (see Fig. 11). Then, it degrades because the increased generating time reduces the available time for communication, leading to a drop in content quality (see Fig. 11). Moreover, the optimal generating step in SAQA-FG increases with computing capacity $f$. For $f$ = 16, the optimal step is 7, whereas for $f$ = 20, it increases to 9. This is because a larger $f$ can support a larger generating step without impacting the communication time (content quality), thereby mitigating the impact of AEG and allowing the assumption of perfect generating accuracy to approximately hold. It can be seen that the system performance is sensitive to the generating step, which is a key factor in balancing content accuracy and quality (see Fig. 11). Thus its optimization is crucial for maximizing AvgCAQA.
@@
6 Conclusion
@@
Considering the imperfect content accuracy in ISAC-based AIGC networks caused by inaccurate sensing input and the inherent AEG during the generating process, this paper proposed a content accuracy and quality aware service metric CAQA. Based on CAQA, a joint sensing, generating, and communication resource allocation problem (CAQA-AIGC) was formulated, aiming to maximize average CAQA (AvgCAQA). To solve this problem with low complexity, the LPDRL-F algorithm was proposed. Compared with existing DRL-based and GDM-based algorithms, LPDRL-F achieves over 14% improvement in AvgCAQA by operating in a smaller and more targeted solution space. Moreover, compared to the CGQ-only oriented scheme, the proposed CAQA-AIGC with LPDRL can improve AvgCAQA by over 50%. This study demonstrates that in practical ISAC-based AIGC networks with imperfect content accuracy caused by inaccurate sensing and AEG, CAQA can serve as an effective performance metric to guide multi-domain resource co-optimization (sensing, computing, and communication), thereby enhancing the overall service quality of generated content. Future work will focus on a more practical ISAC-based AIGC networks with collaborative sensing among multiple ISAC devices and coordinated generation across multiple AIGC-enabled edge computing nodes.
Appendix A Pseudocode, Optimality Proof, and Complexity for RCE Algorithm

### v2

@@ IV-D Complexity Analysis
For the overall complexity of LPDRL-F, if standard LP solvers, such as simplex methods, are employed to solve ComRA sub-problem, the complexity of simplex methods cannot be ignored, and the overall complexity of LPDRL-F would be $O\left({{2^{K}}+\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}}\right)$. In contrast, the time complexity of our proposed LPDRL-F algorithm can be summarized as $O\left({K\log K+\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}}\right)$. Furthermore, the complexity can be further reduced to $O\left({\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}}\right)$ in typical immersive VR/XR interaction scenarios. This is because existing DRL typically use 32-256 hidden layer neurons [38]. According to 3GPP standards [39, 40], the number of immersive VR/XR interaction users $K$ is at the order of 10. Thus, $K\log K\ll\sum\nolimits_{l=1}^{L}{{q_{l}}{q_{l+1}}}$, and the complexity of the RCE algorithm can be neglected.
V Performance Evaluation
@@
Minimum CAQA $\Omega_{k,\text{min}}$ $\sim U[0.4,0.45]$
@@
Simulation parameters of the ISAC-based AIGC environment and LPDRL-F are listed in Table I and Table II, respectively [17, 18, 20, 21, 41, 42]. Large-scale path loss is given by $128.1+37.6{\rm{lo}}{{\rm{g}}_{10}}\left({d/1000}\right)$, where $d$ is the distance between the ISAC device and the user in meters [33]. $\delta^{2}$} is -174 dBm/Hz. The simulation consists of 6000 training episodes, each with 50 iterations, and 10000 testing episodes with 10 iterations. To fully evaluate the algorithm with diverse scenarios, user positions are randomly reset at the beginning of each episode, and users’ task priority level and CAQA requirements are reset at each time slot. Unless otherwise stated, all simulations are conducted with the above settings.
@@
  • DRL-F with Uniform communication (UDRL-F): Sensing and generating resource allocation (i.e., optimizing sensing energy and generating step) are solved using DRL-F, and the remaining energy is evenly allocated to all users for communication after meeting the minimum CAQA requirement.
@@
  • Sensing accuracy and quality aware with fixed generating step (SAQA-FG): Consider the impact of sensing accuracy and communication quality while assuming perfect generation without error. i.e., AEG=0 [20]. Sensing energy allocation is solved using DRL-F with an action filter, and ComRA is optimized by LP. Since AEG is not considered, the generating steps are fixed to a given value. Comparing SAQA-FG with CAQA-AIGC can show the importance of considering AEG in ISAC-AIGC networks.
@@
It can be concluded that the proposed LPDRL-F achieves better convergence and higher AvgCAQA than the generative diffusion model (GDM) and DRL baselines. Therefore, the superiority of the proposed LPDRL-F over SOTA algorithms such as GDM and RL is verified. Then, in Figs. 611, we tried to verify the effectiveness and importance of introducing the content accuracy in the service assessment metric, i.e., the proposed content accuracy and quality aware service assessment metric (CAQA). Therefore, these figures compare the proposed CAQA-AIGC scheme (with the proposed LPDRL-F algorithm) with existing schemes that consider only content generation quality (CGQ) (with a CGQ-FSG-$\alpha$ algorithm [44]) or both sensing accuracy and content generation quality (SAQA) (with a SAQA-FG algorithm [20]).
@@
Then, Fig. 7 shows the AvgCAQA performance as a function of the maximum total energy $E_{\max}$}. It can be seen that for all investigated algorithms, AvgCAQA improves as $E_{\max}$} increases, due to more available energy for both sensing and communication. As $E_{\max}$} continues to increase, the gain in AvgCAQA gradually saturates. This is because AvgCAQA is jointly constrained by both maximum energy and service time (as shown in C1–C3 in (9)), which are coupled with each other. Note that with a fixed maximum service time, the energy that can be utilized is inherently capped. Thus, as the energy continues to increase and reaches this usable limit, further increasing the energy no longer yields noticeable improvement. As shown in Fig. 12, further AvgCAQA gains may be achieved by increasing the maximum service time. The proposed CAQA-AIGC with LPDRL-F consistently achieves the highest AvgCAQA. Similar to Fig. 6, this is because with the assumption of perfect content accuracy, CGQ-FSG-$\alpha$ is designed to ensure a fixed sensing energy proportion and then allocate the remaining energy for communication based on a CGQ-oriented strategy. In a practical system with imperfect content accuracy, when the total energy is small, allocating a larger proportion of sensing energy can achieve higher content accuracy (see Fig. 8). However, it also results in lower content quality since less resource is available for communications (see Fig. 9). Compared to CGQ-FSG-$\alpha$, CAQA-AIGC with LPDRL-F can achieve a better balance between the content accuracy and quality and outperforms CGQ-FSG-$\alpha$ when $E_{\max}$} is small. As $E_{\max}$} increases, CGQ-FSG-$\alpha$ can allocate sufficient energy for communication and enhance content quality (see Fig. 9). However, since a fixed sensing energy is employed by CGQ-FSG-$\alpha$, it cannot fully exploit the advantage of high total energy to further enhance content accuracy. Different to CGQ-FSG-$\alpha$, CAQA-AIGC is aware of the content accuracy in a practical system. It can optimize the sensing-generating-communication resource allocation to fully exploit the high total energy. As shown in Fig. 8 and Fig. 9, when $E_{\max}$} is large, CAQA-AIGC achieves a notably better content accuracy with a slightly worse content quality. Overall, the AvgCAQA performance of CAQA-AIGC is always better than that of CGQ-FSG-$\alpha$.
@@
To evaluate the impact of AEG (one of the sources affecting content accuracy) and the effectiveness of adaptive generating step optimization, Fig. 10 shows AvgCAQA as a function of the generating steps with different computing capacities. SAQA-FG uses a fixed number of generating steps while optimizing sensing and communication energy. As the generating steps increase, its AvgCAQA first improves due to higher generation accuracy (see Fig. 11), then degrades because longer generation time reduces the time available for communication, which lowers content quality (see Fig. 11). It should be noted that with CAQA-AIGC, instead of using fixed generating steps, the generating steps are jointly OPTIMIZED with sensing and communication resources, yielding consistently high AvgCAQA with various computing capacities. Moreover, the optimal generating step in SAQA-FG increases with computing capacity $f$. For $f$ = 16, the optimal step is 7, whereas for $f$ = 20, it increases to 9. This is because a larger $f$ can support a larger generating step without impacting the communication time (content quality), thereby mitigating the impact of AEG and allowing the assumption of perfect generating accuracy to approximately hold. It can be seen that the system performance is sensitive to the generating step, which is a key factor in balancing content accuracy and quality (see Fig. 11). Thus its optimization is crucial for maximizing AvgCAQA.
@@
Figure 13: Convergence of LPDRL-F with different numbers of users.
Note that according to 3GPP standard [39, 40], typically around ten VR users are assumed. When larger user scales are considered, more resources are needed. So Fig. 13 and Fig. 14 are obtained with the settings of ${E_{\max}}=4.0$, ${T_{\max}}=4.0$, $B=500{\mathop{\rm MHz}\nolimits}$, and $2\times{10^{4}}$ training episodes. Fig. 13 shows the convergence performance of the proposed LPDRL-F with larger user scales, where the lines represent the average AvgCAQA and the shaded regions indicate the standard deviation of AvgCAQA. A smaller shaded area stands for better training stability and lower performance variation. As the number of episodes increases, the standard deviation decreases. This is because the policy of LPDRL-F becomes increasingly stable as interactions with the environment increase. As the number of users $K$ increases, the shaded area increases, and so do the training episodes (training time) required to reach convergence. This is because a larger $K$ introduces stronger inter-user coupling in the sensing, generation, and communication resource allocation. The dimension of decision variables is $2K$, resulting in a more complex state-action mapping and increased learning difficulty, which leads to higher performance variability in the early stages and slows the convergence.
Figure 14: AvgCAQA versus the number of users.
Fig. 14 shows the performance of the proposed LPDRL-F and the comparing algorithms as a function of the number of users $K$. It can be seen that the performance of all concerned algorithms decreases as $K$ increases, because resource contention intensifies. The proposed LPDRL-F achieves the best AvgCAQA performance. The AvgCAQA of LPDRL without the action filter drops sharply, indicating that the filter is essential for removing infeasible actions, especially at large $K$. This is because, as $K$ increases, resource contention becomes more intense and the proportion of infeasible actions grows, making the action filter increasingly effective. Meanwhile, the AvgCAQA of UDRL-F, which allocates communication resources uniformly, is worse than that of LPDRL-F, which yields optimal communication resource allocations with the help of LP guidance. Overall, the results show that LP guidance and the action filter are both necessary to sustain high performance. When $K$ is small ($K\leq 30$), LPDRL outperforms UDRL-F because resources are relatively sufficient, and the LP module can exploit communication resources more effectively than uniform allocation, making the LP more effective.
VI Conclusion and Discussion
@@
Considering the imperfect content accuracy in ISAC-based AIGC networks caused by inaccurate sensing input and the inherent AEG during the generating process, this paper proposed a content accuracy and quality aware service metric CAQA. Based on CAQA, a joint sensing, generating, and communication resource allocation problem (CAQA-AIGC) was formulated, aiming to maximize average CAQA (AvgCAQA). To solve this problem with low complexity, the LPDRL-F algorithm was proposed. Compared with existing DRL-based and GDM-based algorithms, LPDRL-F achieves over 10% improvement in AvgCAQA by operating in a smaller and more targeted solution space. Moreover, compared to the CGQ-only oriented scheme, the proposed CAQA-AIGC with LPDRL can improve AvgCAQA by over 50%. This study demonstrates that in practical ISAC-based AIGC networks with imperfect content accuracy caused by inaccurate sensing and AEG, CAQA can serve as an effective performance metric to guide multi-domain resource co-optimization (sensing, computing, and communication), thereby enhancing the overall service quality of generated content.
The overhead of the LPDRL-F agent’s state collection and resource-decision conveying is critical [45, 46]. First, the entire LPDRL-F agent, including the DRL-F module for SGenRA and the RCE module for ComRA, is executed on the edge ISAC device. Considering the overhead of state collection for the agent, since the agent only requires each user’s channel gain and a single scalar-valued CAQA requirement, this information can be reported by each user to the agent with an uplink transmission of tens of bytes, resulting in low communication overhead [47]. For resource decision conveying, since the sensing and communication allocations are executed at the edge ISAC device, no communication overhead is incurred. The AIGC generation-step decisions are executed at the AIGC server. This requires transmitting only $K$ scalar values (one per user) over the backhaul between the ISAC device and the AIGC server, so this overhead is negligible compared to the transmission of raw AIGC data [48, 49]. Since this paper focuses on resource allocation, a subjective validation is beyond the scope of this paper. Such validation may be incorporated in future work through user studies or perceptual experiments to further strengthen the physical meaning of CAQA and enhance its alignment with human perception. It should be noted that, as an initial work in resource allocation for ISAC-driven AIGC networks, this paper considers a basic scenario with one single ISAC device and one AIGC server. However, in practice, a single device and an AIGC server may suffer from limited processing capacity. Future work should focus on more practical ISAC-based AIGC networks with multiple ISAC devices and AIGC servers, where collaborative sensing and coordinated generation across AIGC-enabled edge computing nodes would be the main challenge.
Appendix Appendix A Pseudocode for LPDRL-F Algorithm
Initialize actor, critic, and target critic parameters $\phi$, $\theta_{j}$, $\theta^{\prime}_{j}$ for $j\in\{1,2\}$, and replay buffer $\Psi$.
for each episode to max episode do
 for iteration $i=1$ to max iteration do
      Collect current state $s_{i}$ and select action $a_{i}$ from the current policy.
     Decode $a_{i}$ to obtain the generating steps $\mathbf{z}(i)$ and the original sensing energy allocation.
     Apply the action filter in (10) to obtain the sensing energy allocation $\mathbf{E}_{s}(i)$.
     Run the RCE algorithm to obtain the communication energy allocation $\mathbf{E}_{c}(i)$.
     Compute reward $r_{i}$ using (11) and observe next state $s_{i+1}$.
     Store transition $\{s_{i},a_{i},r_{i},s_{i+1}\}$ into $\Psi$.
    if $|\Psi|>\Psi_{\min}$ then
         Sample a batch $\Psi_{b}$ from $\Psi$.
        Update the critic network via (13).
        Update the actor network via (14).
        Update the target critic network via (15).
        Update the temperature parameter $\rho$ via $\nabla_{\rho}J(\rho)$ [31].
Algorithm 1 LPDRL-F for CAQA-AIGC Problem
Appendix Appendix B Pseudocode, Optimality Proof, and Complexity for RCE Algorithm